A 2kg puck moves to the right with an initial speed of 5 m/s and collided with a stationary 3kg puck. Assuming an elastic collision (kinetic energy is conserved) find the final velocity of each puck.
1
Expert's answer
2019-03-06T09:53:40-0500

Let's write the law of conservation of momentum:


"m_1v_{1i} = m_1v_{1f} + m_2v_{2f}, (1)"

here, "m_1", "m_2" are the masses of the 2-kg and 3-kg pucks, respsectively; "v_{1i}" is the initial velocity of the 2-kg puck which moves to the right; "v_{1f}", "v_{2f}" are the final velocities of the 2-kg and 3-kg pucks, respectively.

Let's write the law of conservation of energy:


"\dfrac{1}{2} m_1v_{1i}^2 = \dfrac{1}{2} m_1v_{1f}^2 + \dfrac{1}{2} m_2v_{2f}^2. (2)"

Let’s transpose the terms with "m_1" to the left-side of the equations (1) and (2), respectively:


"m_1(v_{1i} - v_{1f}) = m_2v_{2f}, (3)""m_1(v_{1i}^2 - v_{1f}^2) = m_2v_{2f}^2. (4)"

Dividing equation (4) by equation (3), we get:


"v_{1i} + v_{1f} = v_{2f}. (6)"


Substituting equation (6) into the equation (3), we get:


"m_1v_{1i} - m_1v_{1f} = m_2v_{1i} + m_2v_{1f},"


From this equation we can find "v_{1f}":

"v_{1f} = \dfrac{m_1 - m_2}{m_1 + m_2} \cdot v_{1i} = \dfrac{2 kg - 3 kg}{2 kg + 3 kg} \cdot 5 \dfrac{m}{s} = -1 \dfrac{m}{s}."


The sign minus indicates that the 2-kg puck after the collision moves in the opposite direction (to the left).

Finally, from the equation (6) we can find "v_{2f}":


"v_{2f} = v_{1i} + v_{1f} = 5 \dfrac{m}{s} + (-1 \dfrac{m}{s}) = 4 \dfrac{m}{s}."

The sign plus indicates that the 3-kg puck after the collision moves to the right.

Answer:

"v_{1f} = -1 \dfrac{m}{s}", "v_{2f} = 4 \dfrac{m}{s}."